# Self-Tuning: Instructing LLMs to Effectively Acquire New Knowledge through Self-Teaching

### v3

@@ Abstract
Large language models (LLMs) often struggle to provide up-to-date information due to their one-time training and the constantly evolving nature of the world. To keep LLMs current, existing approaches typically involve continued pre-training on new documents. However, they frequently face difficulties in extracting stored knowledge. Motivated by the remarkable success of the Feynman Technique in efficient human learning, we introduce Self-Tuning, a learning framework aimed at improving an LLM’s ability to effectively acquire new knowledge from raw documents through self-teaching. Specifically, we develop a Self-Teaching strategy that augments the documents with a set of knowledge-intensive tasks created in a self-supervised manner, focusing on three crucial aspects: memorization, comprehension, and self-reflection. In addition, we introduce three Wiki-Newpages-2023-QA datasets to facilitate an in-depth analysis of an LLM’s knowledge acquisition ability concerning memorization, extraction, and reasoning. Extensive experimental results on Llama2 family models reveal that Self-Tuning consistently exhibits superior performance across all knowledge acquisition tasks and excels in preserving previous knowledge.††∗Work done during the internship at Tencent AI Lab. 111Our source code and the curated new datasets can be accessed publicly at https://github.com/zhangxy-2019/Effective-Knowledge-Injection.
@@ 1 Introduction
Armed with a wealth of factual knowledge acquired during the pre-training phase Zhou et al. (2023), LLMs Touvron et al. (2023); OpenAI (2023) exhibit remarkable proficiency in numerous knowledge-intensive tasks Zhang et al. (2024c); Cohen et al. (2023); Gekhman et al. (2024). Despite this, the knowledge stored in LLMs can quickly become outdated due to the one-time training of LLMs and the ever-changing nature of the world Huang et al. (2023); Jiang et al. (2024b). These unavoidable knowledge limitations present notable obstacles to the trustworthiness of LLMs in real-world scenarios Liu et al. (2023); Mecklenburg et al. (2024). Therefore, it is essential to equip LLMs with new knowledge to keep them up-to-date.
In this paper, we focus on injecting new knowledge into the parameters of LLMs. As depicted in the upper part of Figure 1, a standard approach involves continued pre-training (A) on a raw corpus (here, test doc) containing new information Jang et al. (2022). However, it struggles to extract the embedded knowledge, potentially due to the impaired question-answering (QA) capability Allen-Zhu and Li (2023); Cheng et al. (2024). Despite the assistance of subsequent instruction-tuning (B) Wei et al. (2022); Ouyang et al. (2022a) on QA data, the knowledge retrieved from the LLMs remains notably constrained Jiang et al. (2024b). Recently, Jiang et al. (2024b) suggests fine-tuning on a mix of QA data and related documents before continuing pre-training, with the aim of teaching the model how to access knowledge from documents and answer questions. Although this method greatly outperforms standard approaches, our initial results suggest that its effectiveness in knowledge extraction remains limited.
@@
In addition, we introduce three Wiki-Newpages-2023-QA datasets to conduct an in-depth study of how an LLM acquires new knowledge w.r.t., memorization, extraction, and comprehension (in this study, reasoning) across single-domain, multi-domain, and cross-domain settings. These datasets are carefully curated to ensure minimal overlap with the LLM’s pre-training corpora, emphasizing two key knowledge-intensive tasks, i.e., open-ended generation and natural language inference (NLI) tasks. Extensive experimental results on Llama2 family models demonstrate that Self-Tuning significantly outperforms all other compared methods on knowledge memorization and extraction tasks. In addition, Self-Tuning consistently yields high accuracy on reasoning tasks, while the performance of the compared methods largely fluctuates in different scenarios. Inspiringly, Self-Tuning exhibits exceptional performance in retaining previously acquired knowledge (i.e., knowledge retention) concerning extraction and reasoning on two well-established benchmarks.
@@
2 Related Work
Continual Knowledge Injection.
The primary research approach for injecting new knowledge into LLMs Xu et al. (2023); Ovadia et al. (2024); Mecklenburg et al. (2024) is through continued pre-training. This method entails the ongoing pre-training of LLMs on raw corpora containing new knowledge, carried out in a causal auto-regressive manner Allen-Zhu and Li (2023); Ibrahim et al. (2024); Ovadia et al. (2024). However, this straightforward approach often encounters hurdles in effectively enabling LLMs to extract the acquired knowledge during the inference phase Allen-Zhu and Li (2023); Jiang et al. (2024b); Cheng et al. (2024). To enhance knowledge extraction, instruction tuning on QA data after pre-training has been extensively employed Wei et al. (2022); Ouyang et al. (2022b). Jiang et al. (2024b) suggests that the effectiveness of this method remains limited, and proposes fine-tuning the model on QA data before continued pre-training. This instructs the model on how to retrieve knowledge from raw corpora, thereby enhancing knowledge extraction. However, such an approach tends to underestimate the importance of comprehending the new knowledge.
Recognizing the significance of knowledge comprehension, Cheng et al. (2024) proposes converting raw corpora into reading comprehension texts. This approach, however, focuses on domain adaptation and preserving general prompting abilities by mining a set of instruction-following tasks from the document content. In contrast, our work aims to equip the model with the ability to effectively absorb new knowledge from raw documents and employ the learned ability to unseen documents. Specifically, we develop a Self-Teaching strategy to present the raw document as plain texts for memorization, accompanied by a set of tasks for comprehension and self-reflection, which are created based on raw corpora in a self-supervised manner, without relying on any mining patterns.
Retrieval-Augmented Generation.
@@ 3 Wiki-Newpages-2023-QA: Datasets for Studying LLM Knowledge Acquisition
| Wiki-Film (Single-domain) | Genre, Language, Director, Released Time, etc. | Test: 955 (# QA); 169 (# Docs) | 8.83 (Q) 4.61 (A) 58.10 (Doc) | 1,387 (# QA) 169 (# Docs) | Yes (62.73%) No (26.53%) Impossible (2.52%) |

### v4

@@ Abstract
Large language models (LLMs) often struggle to provide up-to-date information due to their one-time training and the constantly evolving nature of the world. To keep LLMs current, existing approaches typically involve continued pre-training on new documents. However, they frequently face difficulties in extracting stored knowledge. Motivated by the remarkable success of the Feynman Technique in efficient human learning, we introduce Self-Tuning, a learning framework aimed at improving an LLM’s ability to effectively acquire new knowledge from unseen raw documents through self-teaching. Specifically, we develop a Self-Teaching strategy that augments the documents with a set of knowledge-intensive tasks created in a self-supervised manner, focusing on three crucial aspects: memorization, comprehension, and self-reflection. Additionally, we introduce three Wiki-Newpages-2023-QA datasets to facilitate an in-depth analysis of an LLM’s knowledge acquisition ability concerning memorization, extraction, and reasoning. Extensive experimental results on various models, e.g., Llama2-7B reveal that Self-Tuning consistently exhibits superior performance across all knowledge acquisition tasks and excels in preserving previous knowledge.
@@ 1 Introduction
Armed with a wealth of factual knowledge acquired during the pre-training phase sections, LLMs Touvron et al. (2023a); OpenAI (2023) exhibit remarkable proficiency in numerous knowledge-intensive tasks Cohen et al. (2023); Zhang et al. (2023); Gekhman et al. (2024). Despite this, the knowledge stored in LLMs can quickly become outdated due to the one-time training of LLMs and the ever-changing nature of the world Huang et al. (2023); Jiang et al. (2024c). These unavoidable knowledge limitations present notable obstacles to the trustworthiness of LLMs in real-world scenarios Liu et al. (2023); Mecklenburg et al. (2024). Thus, it is essential to equip LLMs with new knowledge to keep them up-to-date.
In this paper, we focus on injecting new knowledge into the parameters of LLMs. As depicted in the upper part of Figure 1, a standard approach involves continued pre-training (A) on a raw corpus (here, test doc) containing new information Jang et al. (2022). However, it struggles to extract the embedded knowledge, potentially due to the impaired question-answering (QA) capability Allen-Zhu and Li (2023); Cheng et al. (2024). Despite the assistance of subsequent instruction-tuning (B) Wei et al. (2022); Ouyang et al. (2022a) on QA data, the knowledge retrieved from the LLMs remains notably constrained Jiang et al. (2024c). Recently, Jiang et al. (2024c) suggests fine-tuning on a mix of QA data and related documents before continuing pre-training, with the aim of teaching the model how to access knowledge from documents and answer questions. Although this method greatly outperforms standard approaches, our initial results suggest that its effectiveness in knowledge extraction remains limited.
@@
In addition, we introduce three Wiki-Newpages-2023-QA datasets to conduct an in-depth study of how an LLM acquires new knowledge w.r.t., memorization, extraction, and comprehension (in this study, reasoning) across single-domain, multi-domain, and cross-domain settings. These datasets are carefully curated to ensure minimal overlap with the LLM’s pre-training corpora, emphasizing two key knowledge-intensive tasks, i.e., open-ended generation and natural language inference (NLI) tasks. Extensive experimental results on diverse models, e.g., Llama2-7B Touvron et al. (2023b), Qwen2-7B Yang et al. (2024), and Mistral-7B-v0.1 Jiang et al. (2023) demonstrate that Self-Tuning significantly outperforms all other compared methods on knowledge memorization and extraction tasks. In addition, Self-Tuning consistently yields high accuracy on reasoning tasks, while the performance of the compared methods largely fluctuates in different scenarios. Inspiringly, Self-Tuning exhibits exceptional performance in retaining previously acquired knowledge (i.e., knowledge retention) concerning extraction and reasoning on two well-established benchmarks.
@@
| Wiki-Film (Single-domain) | Genre, Language, Director, Released Time, etc. | Test: 955 (# QA); 169 (# Docs) | 8.83 (Q) 4.61 (A) 58.10 (Doc) | 1,387 (# QA) 169 (# Docs) | Yes (62.73%) No (26.53%) Impossible (10.74%) |
2 Related Work
Continual Knowledge Injection.
The primary research approach for injecting new knowledge into LLMs Xu et al. (2023); Ovadia et al. (2024); Mecklenburg et al. (2024) is through continued pre-training. This method entails the ongoing pre-training of LLMs on raw corpora containing new knowledge, carried out in a causal auto-regressive manner Allen-Zhu and Li (2023); Ibrahim et al. (2024); Ovadia et al. (2024). However, this straightforward approach often encounters hurdles in effectively enabling LLMs to extract the acquired knowledge during the inference phase Allen-Zhu and Li (2023); Jiang et al. (2024c); Cheng et al. (2024). To enhance knowledge extraction, instruction tuning on QA data after pre-training has been extensively employed Wei et al. (2022); Ouyang et al. (2022b). Jiang et al. (2024c) suggests that the effectiveness of this method remains limited, and proposes fine-tuning the model on QA data before continued pre-training. This instructs the model on how to retrieve knowledge from raw corpora, thereby enhancing knowledge extraction. However, such an approach tends to underestimate the importance of comprehending the new knowledge.
Acknowledging the value of knowledge comprehension, Cheng et al. (2024) proposes converting raw corpora into reading comprehension texts. This approach, however, focuses on domain adaptation and preserving general prompting abilities by mining a set of instruction-following tasks from the document content. In contrast, our work aims to equip the model with the ability to effectively absorb new knowledge from raw documents and employ the learned ability to unseen documents. Specifically, we develop a Self-Teaching strategy to present the raw document as plain texts for memorization, accompanied by a set of tasks for comprehension and self-reflection, which are created based on raw corpora in a self-supervised manner, without relying on any mining patterns.
Additionally, knowledge editing Mitchell et al. (2022); Zheng et al. (2023); Yao et al. (2023); Jiang et al. (2024b); Zhang et al. (2024a) and retrieval-augmented generation Lewis et al. (2021); Vu et al. (2023); Ovadia et al. (2024); Jeong et al. (2024) are recognized as two related research fields. We provide more details on these initiatives in Appendix A.
To explore the knowledge acquisition capabilities of LLMs from new documents, w.r.t., memorization, extraction and reasoning, we introduce the Wiki-Newpages-2023-QA datasets (Table 1), which are carefully designed to minimize overlap with the initial pre-training corpus. These datasets comprise new document corpora for studying knowledge memorization and associated QA datasets for two vital knowledge-intensive tasks: open-ended generation and NLI for examining extraction and reasoning, respectively. Due to space constraints, we provide a brief overview of the dataset construction process here, with the complete version available in Appendix B.
3.1 Document Collection and QA Pair Generation
Document Collection.
To construct the document corpus, we collect articles from September to October 2023 (4,257 articles in total) from Wikipedia NewPages111https://en.wikipedia.org/wiki/Special:NewPages, which include new articles from various domains published after the pre-training cut-off time of the LLMs being evaluated.222The pre-training cut-off time for the Llama2 family models used in this study is 2022. Following Jiang et al. (2024c), we only use the first paragraph of each article, as it offers a comprehensive summary and contains a wealth of factual information.
QA Pair Generation.
3.2 Splitting
We construct three datasets for single-domain, multi-domain, and cross-domain analysis, splitting them into training and testing subsets while ensuring zero knowledge overlap.
Dataset Splitting.
We generate three datasets: Wiki-Newpages-2023-10-Bio (Wiki-Bio), Wiki-Newpages-2023-10-Multi (Wiki-Multi), and Wiki-Newpages-2023-(9)10-Film (Wiki-Film) by randomly selecting 1,263 biographical documents, 2,104 multi-domain documents, and 955 film documents from the collected document corpus and their associated QA pairs.
Train-test Splitting.
We divide Wiki-Bio and Wiki-Multi datasets into training and testing subsets for single-domain and multi-domain evaluations. We use Wiki-Film as the test set for cross-domain scenarios. Note that the training QA datasets only include open-ended generation task pairs, ensuring fair comparisons.
4 Self-Tuning
4.1 Overview
Our objective is to equip an LLM , parameterized by , with the ability to learn how to derive knowledge from raw documents. This is achieved by training the model using a combination of training document dataset and associated training QA dataset , as depicted in the upper left part of Figure 2. To enhance effective knowledge absorption, we present along with a series of knowledge-intensive tasks (a.k.a. self-teaching tasks) that are related to their content using the proposed Self-Teaching strategy (in the lower part of Figure 2). These tasks are generated in a self-supervised manner based on the contents of , using the proposed Self-Teaching learning approach (Section 4.2). The multi-task training objective is:
Stage 2: Learn New Knowledge while Reviewing QA Skills.
Our aim is to train the model to apply the learned strategy for spontaneously extracting new knowledge from unseen documents (i.e., raw test corpora ). In addition to training on , we include , allowing the model to review and refine its question-answering ability. The objective of this stage is:
Stage 3: Continually Learn.
Our goal is to ensure that the model thoroughly absorbs the new knowledge by conducting follow-up training on (raw corpora). The objective is as follows:
4.2 Self-Teaching Learning Strategy
Memorization.
Comprehension.
Gist identification improves the model’s ability to pinpoint the key elements (i.e., entities) within the atomic facts. Specifically, we prompt the model with Highlight the key information within the article:, and use the entities within the document as gold answers, identified using Spacy333https://spacy.io/usage.
Natural language inference provides the model with the capability to determine whether a statement can be inferred from specific document contents (i.e., “Yes,” “No,” or “It’s impossible to say”), thus avoiding misconceptions that may arise during knowledge acquisition. Specifically, we use a randomly sampled sentence (identified using NLTK444A natural language toolkit. https://www.nltk.org/) within the document content as the true statement, and a corrupted version where one entity is replaced by an irrelevant entity from another sentence as the false statement. Then, we prompt the model with Based on the article above can we conclude that and the sampled sentence (either initial or corrupted), with the three relations as options and corresponding answers.
Self-Reflection.
5 Experiments
| Method | Training Data in Each Stage | ||
| Stage 1 | Stage 2 | Stage 3 | |
| Continued Pre-Training | test doc | ||
| Standard Instruction-Tuning | train doc&test doc | train QA | |
| PIT | train QAtrain doc | test doc | |
| Variants of Self-Tuning | |||
| Self-Tuning via Read. | train QA&train doc (reading-comprehension format Cheng et al. (2024)) | test doc | |
| Method | Wiki-Newpages-2023-QA (Acquisition) | NQ (Reten.) | CSQA (Reten.) | |||||||
| Memorization | Extraction | Reason. | Extraction | Reasoning | ||||||
| Knowledge Acquisition on Wiki-Newpages-2023-10-Bio (Single-Domain Scenario) | ||||||||||
| w/o Knowledge Injection | ||||||||||
| w/ Knowledge Injection | ||||||||||
| Cont. Pre-training | 7.28 | 6.33 | 3.62 | 15.96 | 18.72 | 16.11 | 15.09 | 16.00 | 24.11 | 53.40 |
| PIT | 1.52 | 6.39 | 4.50 | 16.97 | 18.92 | 17.10 | 3.03 | 13.06 | 23.42 | 54.38 |
5.1 Setup
Datasets and Evaluation Metrics. We validate Self-Tuning in both knowledge acquisition and retention for a well-rounded analysis.
We perform assessments on three knowledge acquisition tasks: Memorization: We use test document datasets and report perplexity (PPL) Jelinek et al. (1977). Extraction: We use test QA datasets for open-ended generation tasks and evaluate factual accuracy using exact match (EM), Recall, F1 Kwiatkowski et al. (2019), Rouge-L Lin (2004), and accuracy. Reasoning: We use test QA datasets for NLI tasks and report accuracy.
We evaluate two aspects of knowledge retention: Extraction: We assess the model’s performance in retaining general factual knowledge using Natural Questions (NQ) Kwiatkowski et al. (2019), with EM and F1. Reasoning: We evaluate the capability in retaining commonsense knowledge using CommonsenseQA (CSQA) Talmor et al. (2019) and report accuracy. All evaluations are conducted in a closed-book setting (see Appendix S).
Compared Methods. We compare Self-Tuning with three representative approaches (Table 2): (1) Continued Pre-training Ovadia et al. (2024), (2) Standard Instruction-tuning Saito et al. (2024), and (3) PIT Jiang et al. (2024c), which trains on and with QA pairs positioned before their corresponding document texts. We also evaluate their variants (Table 9). Results, averaged over three runs, show significant differences in means ( < 0.001), with details in Appendix T.
5.2 Main Results
Table 3 (top) presents the evaluation results on Llama2-7B in relation to knowledge acquisition and retention in the single-domain scenario using the Wiki-Bio dataset.
The curated dataset exhibits minimal overlap with the pre-training data of the LLMs. The extremely low performance in the closed-book setting (e.g., with EM around 2% for knowledge extraction) indicates that the dataset has little in common with the pre-training data, thus ensuring the reliability of the evaluation results. The non-zero EM values might be due to a small number of collected Wikipedia articles that were initially published but underwent revisions after the cut-off time.
Self-Tuning substantially improves the LLM’s knowledge acquisition ability. Self-Tuning greatly enhances the performance of Llama2-7B across three dimensions: reducing PPL to nearly 1, signifying effective memorization of the new documents; increasing EM by roughly 11.5% on the knowledge extraction task, attaining performance comparable to the open-book setting; achieving high accuracy among the compared methods for the reasoning task, demonstrating excellent understanding of the newly acquired knowledge. These results underscore the importance of first training the model to develop knowledge absorption capabilities before exposing it to test documents, aligning with findings from Jiang et al. (2024c). More importantly, Self-Tuning significantly outperforms PIT, highlighting the role of comprehension and self-reflection beyond mere memorization, further validating the effectiveness of Self-Teaching.
Self-Tuning excels in knowledge retention. Unlike other methods that display fluctuating performance, Self-Tuning shows a strong ability to maintain previously acquired knowledge in terms of both knowledge extraction and reasoning. The slight improvements in evaluation metrics, such as F1 (roughly 1% on extracting learned world knowledge) and accuracy (around 13% on commonsense reasoning), compared to the closed-book performance without knowledge injection, suggest that systematically learning new knowledge doesn’t necessarily lead to catastrophic forgetting.
We further validate the efficacy of Self-Tuning through additional analyses presented in the following appendices: Appendix C (evaluation of three additional methods to validate the effectiveness of our three-stage design), N (comparison of training on test documents with accompanying QA), D (fine-grained evaluation of the open-ended generation task), E (qualitative comparison of responses from PIT and Self-Tuning), F (ablation study investigating the impact of removing comprehension and self-reflection tasks from self-teaching on knowledge memorization and acquisition), G (error analysis), and K (assessment of training efficiency).
5.3 Results in the Multi-Domain and Cross-Domain Scenarios
To explore the potential of Self-Tuning for enhancing LLM’s knowledge acquisition and retention in real-world scenarios, we evaluate its performance in two challenging settings (Table 3): the multi-domain scenario (in the middle part); the cross-domain scenario (in the bottom part), where the training data is from Wiki-Bio, while the test data is from Wiki-Film.
Self-Tuning shows strong potential in enhancing knowledge acquisition and retention across documents containing diverse new knowledge. In Table 3, Self-Tuning consistently achieves the best performance in both settings.
The capacity to systematically absorb knowledge improves generalization ability. The substantial improvements over all compared methods in the cross-domain setting, e.g., exceeding EM by 13% on the knowledge extraction task, highlight the value of equipping the model with the ability to effectively absorb knowledge from raw documents using the Self-Teaching strategy, rather than solely teaching it how to answer questions.
5.4 Training Dynamics
The remarkable performance of Self-Tuning in enhancing knowledge acquisition does not stem from early-fitting. In Figure 3, we observe that Self-Tuning not only memorizes new knowledge more rapidly than the compared methods, lowering PPL to almost 1 within 3 epochs, but also consistently achieves the best performance during long-term training. Remarkably, Self-Tuning begins to outperform the open-book performance from the 5th epoch and reaches its peak at the 25th epoch with a 5% higher EM score on the knowledge extraction task.
Self-Tuning performs well in preserving previously acquired knowledge, with only a small decline in EM of roughly 2-3% over the course of 50 training epochs. This suggests that Self-Tuning has great potential for real-world applications.
5.5 Variants of Self-Tuning
| Method | Wiki-Bio (Acquisition) | ||||
| Mem. | Extraction | Reason. | |||
| Cont. Pre-training | 7.28 | 4.68 | 2.87 | 15.07 | 7.96 |
[…]
| S.T. via Read. | 1.46 | 20.97 | 17.65 | 34.55 | 39.37 |
| S.T. w/ Pre-Review | 1.28 | 29.86 | 25.94 | 43.31 | 46.91 |
| Self-Tuning | 1.11 | 37.25 | 31.52 | 50.61 | 44.31 |
Setup. We investigate three variants of Self-Tuning (in the lower part of Table 2): Self-Tuning w/o Review, where we continue training on test documents without the reviewing capability; Self-Tuning via Read., which displays the training documents in a reading-comprehension format Cheng et al. (2024) (see Table 28); Self-Tuning w/ Pre-Review, which trains on a mix of training documents and training QA in the second stage, before training on test documents.
Results. In Table 5.5, despite having lower performance than Self-Tuning, all variants significantly enhance the model’s ability for knowledge acquisition compared to continued pre-training.
Reviewing QA ability aids in knowledge acquisition. Compared to Self-Tuning, Self-Tuning w/o Review exhibits inferior performance. Moreover, we suspect that the lower performance of Self-Tuning w/ Pre-Review is because reviewing QA ability during, rather than before, the continuous learning of new knowledge is more effective in reducing distribution shift, thereby stabilizing the training process.
5.6 Results of Varying Models and Corpora
| Method | Acquisition | ||||
| Varying Model (Qwen2-7B on WikiBio-2024) | |||||
| Closed-book | 12.41 | 4.16 | 2.55 | 15.01 | 13.17 |
| Stand. Ins.-tuning | 2.77 | 11.29 | 9.36 | 25.45 | 24.83 |
| PIT | 1.97 | 11.41 | 9.53 | 25.98 | 25.64 |
| Self-Tuning | 1.14 | 31.79 | 28.51 | 44.91 | 43.33 |
| Varying Model (Mistral-7B-v0.1 on WikiBio-2023) | |||||
| Closed-book | 8.45 | 6.64 | 4.37 | 19.51 | 17.25 |
| Stand. Ins.-tuning | 2.84 | 16.44 | 13.88 | 29.54 | 29.13 |
| PIT | 1.42 | 26.85 | 23.08 | 40.36 | 39.52 |
| Self-Tuning | 1.08 | 41.63 | 36.50 | 55.32 | 52.87 |
| Varying Corpora (Llama2-7B on WebNews-2023) | |||||
| Closed-book | 11.20 | 9.04 | 6.30 | 24.22 | 17.99 |
| Stand. Ins.-tuning | 3.27 | 21.48 | 13.38 | 37.66 | 31.31 |
| PIT | 1.67 | 30.37 | 18.96 | 51.17 | 40.53 |
| Self-Tuning | 1.10 | 37.48 | 28.74 | 56.26 | 48.21 |
Setup. We evaluate Self-Tuning using diverse models, including Qwen2-7B Yang et al. (2024), Mistral-7B-v0.1 Jiang et al. (2023), and Gemma-7B Team et al. (2024) (see Appendix L), as well as different corpora, such as WebNews-2023 Tang and Yang (2024), which consists of worldwide news articles from diverse websites (see Appendix U for further details).
Results. The results in Table 5 demonstrate that Self-Tuning consistently achieves the best performance, highlighting its strong generalizability across both models and corpora. Further evaluation results for Llama2-13B and Llama2-7B-chat are available in Appendices I and J, respectively.
6 Conclusion
In this study, we introduce Self-Tuning to enhance an LLM’s ability to effectively learn from raw documents through self-teaching. Specifically, we develop Self-Teaching, a self-supervised learning strategy that presents documents as plain texts along with various knowledge-intensive tasks derived directly from the documents. Additionally, we present three Wikipedia-Newpages-2023-QA datasets to enable a comprehensive evaluation of an LLM’s knowledge acquisition capabilities across three distinct scenarios. Our findings show that Self-Tuning consistently yields superior performance on the knowledge acquisition tasks while showing impressive knowledge retention performance. These results suggest the potential for broader applications of Self-Tuning.
Limitations
Combining with Continual Learning Approaches.
Our study primarily focuses on enhancing a language model’s ability to effectively learn new knowledge from previously unseen raw corpora. Although experimental results on MCQA and NQ demonstrate that our Self-Tuning method well preserves previously acquired knowledge, future research could explore integrating Self-Tuning with continual learning approaches Wang et al. (2024). For instance, regularization-based methods such as EWC Kirkpatrick et al. (2017) and replay-based methods, like incorporating segments from general domain datasets (e.g., Wiki data Zhang et al. (2024c)), could improve the model’s capacity to retain learned knowledge and skills while mitigating the risk of overfitting to new information.
In this study, we intentionally avoided using continual learning approaches to ensure a fair comparison of knowledge injection with previous methods. However, we present preliminary results of combining Self-Tuning with a replay-based approach in Appendix M. These results confirm the strong potential of integrating Self-Tuning with continual learning techniques to improve both knowledge acquisition and retention.
Regarding Resource Demands.
To verify the efficacy of Self-Tuning, we provide a detailed analysis of training efficiency on the Wiki-Newpages-2023-Bio dataset, conducted using 8 Tesla V100 GPUs (32GB) with Llama-7B, in Appendix K. This analysis demonstrates that our Self-Tuning framework not only significantly outperforms the strongest baseline method but also achieves this with reduced training time. Furthermore, the effectiveness of Self-Tuning across three distinct scenarios highlights its ability to directly assimilate new knowledge from incoming test documents without requiring retraining on the original training corpus (i.e., omitting the first stage). Notably, Self-Tuning eliminates the need for any additional annotation costs. All experiments were conducted on 8 Tesla V100 GPUs (32GB), with training completing in just a few hours. Consequently, we anticipate minimal barriers to the adoption of Self-Tuning, even for teams with limited computational resources.
Ethics Statement
Throughout our research, we have consistently adhered to ethical guidelines to uphold privacy, fairness, and the well-being of all individuals and groups involved. All benchmark datasets utilized in this study are used solely for research purposes and do not contain any personally identifiable information, thereby safeguarding privacy. During the QA data collection process for Wiki-Newpages-2023-QA, WikiBio-2024, and WebNews-2023 using GPT-4, we meticulously crafted prompts to eliminate any language that might discriminate against specific individuals or groups. These measures were implemented to minimize potential negative effects on users’ well-being. Examples of these carefully designed prompts can be found in Table 23, Table 24, and Table 26. To further ensure the quality of the newly collected datasets, the authors manually reviewed them following the guidelines in Bai et al. (2022). These datasets were confirmed to be of high quality, free from offensive content, false information, and any personally identifiable information Radharapu et al. (2023); Zhou et al. (2023). Future research efforts could explore the OpenAI moderation API555https://platform.openai.com/docs/guides/moderation/overview to systematically filter out inappropriate system responses. Additionally, GPT-4 is solely used to convert document knowledge into QA pairs—without introducing any additional information—and can be replaced by any other model.
Appendix A Additional Efforts for Knowledge Injection
Knowledge editing Zheng et al. (2023); Yao et al. (2023); Jiang et al. (2024b); Zhang et al. (2024a) and retrieval-augmented generation (RAG) Lewis et al. (2021); Ovadia et al. (2024); Jeong et al. (2024) are recognized as two related research initiatives in the field of knowledge injection.
Knowledge editing Mitchell et al. (2022); Zheng et al. (2023); Yao et al. (2023); Jiang et al. (2024b); Zhang et al. (2024a) concentrates on rectifying outdated or inaccurate factual knowledge stored in the model, without affecting other facts. In contrast, our focus lies in enabling LLMs to efficiently acquire knowledge from raw corpora.
Appendix B Wiki-Newpages-2023-QA: Datasets for Studying LLM Knowledge Acquisition